Martinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 2. For financial accounting purposes, what is the total amount of period costs incurred to sell 10,000 units? (Do not round intermediate calculations.) Total period costMartinez Companys relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expensee Fixed administrative expense Sales commissions Variable administrative expense Average Cost Per Unit $5.60 $3.10 $1.40 $4.00 $2.60 $2.20 $1.20 $0.45 3. If 8,000 units are produced and sold, what is the variable cost per unit produced and sold? (Round your answer to 2 decimal places.) Variable cost per unit sold
Answer #1

1.) In the given case, fixed selling expense and sales commission are the period costs incurred to sell the units.

So per unit period costs to sell = fixed selling expense per unit + sales commission per unit = 2.60 + 1.20 = $ 3.80

So period costs to sell 10,000 such units will be 3.80*10,000 = $ 38,000.

2.) Variable costs for production and selling in the case are Direct Material, Direct Labor, Variable Manufacturing Overhead, Sales Commission and Variable administrative overhead. Since production level is 8,000 units, therefore average cost per unit given in the question will be applicable.

So, total variable cost per unit produced and sold, if 8,000 units are produced and sold will be total of average cost per unit for Direct Material, Direct Labor, Variable Manufacturing Overhead, Sales Commission and Variable administrative overhead, i.e., 5.60 + 3.10 + 1.40 + 1.20 + .45 = $ 11.75.